OpenStudy (anonymous):

x=-1,x=2 evaluate the exponential function

OpenStudy (anonymous):

e^-1 and e^2. do you want approximations?

OpenStudy (anonymous):

i am not sure mth is so confusing to me I am barely hanging on.....

jhonyy9 (jhonyy9):

- so if x_1 = -1 and x_2 = 2 you know that the base equation is the form ax*2 +bx +c =0 - and you know that x_1 + x_2 = -b/a and (x_1)(x_2) = c/a - from this you can calculate the a,band c and after you can writing this equation what are those solutions

jhonyy9 (jhonyy9):

- this equation will be (x-1)(x+2)=x*2 +x -2